### v1

# Estimates on the domain of validity for Lyapunov-Schmidt reduction

Abstract
Lyapunov-Schmidt reduction is a dimensionality reduction technique in nonlinear systems analysis that is commonly utilised in the study of bifurcation problems in high-dimensional systems. The method is a systematic procedure for reducing the dimensionality of systems of algebraic equations that have singular points, preserving essential features of their solution sets. In this article, we establish estimates for the region of validity of the reduction by applying bounds on the implicit function theorem derived in [jindal2023imft]. We then apply these bounds to an illustrative example of a two-dimensional system with a pitchfork bifurcation.
I Introduction
A bifurcation in a parameterised nonlinear dynamical system marks a point at which the system undergoes qualitative changes in behaviour, often leading to the creation of multiple new isolated equilibria, onset of oscillation, and chaotic dynamics [guckenheimer2013nonlinear]. Analysis of bifurcations plays an important role in both understanding and controlling systems with complex nonlinear features including power systems [ajjarapu1992bifurcation], epidemic processes [pare2020modeling], social dynamics [bizyaeva2022nonlinear, bizyaeva2023thesis], and neuromorphic computing networks [juarez2024analysis]. In a control context, objectives range from understanding when feedback laws give rise to bifurcations and potentially suppressing their onset [chen1994overview, chen2000bifurcation], to designing dynamics and feedback laws that give rise to meaningful behaviour through a bifurcation [leonard2024fast].
Dimensionality reduction is a key step in formally classifying a bifurcation. While a system of interest may have high state dimension, bifurcations typically occur along low-dimensional manifolds. Identifying low-dimensional representations of systems that preserve their essential features on or near these manifolds allows for a more tractable investigation of dynamics near bifurcation points. Centre manifold reduction and Lyapunov-Schmidt reduction are the two methods most commonly utilised for this purpose.
The centre manifold reduction method is used to simultaneously approximate a nonlinear manifold of a bifurcation along with a low-dimensional representation of the dynamical system with respect to coordinates along this manifold [guckenheimer2013nonlinear, Chapter 3.2]. While this technique is powerful, it requires performing calculations that are challenging and in practice intractable for systems with high state dimension and complex interconnections such as large-scale networks. For such systems, Lyapunov-Schmidt reduction is often a more appropriate tool [sidorov2013lyapunov, golubitsky2023dynamics].
Lyapunov-Schmidt reduction is distinguished from centre manifold reduction by its topological viewpoint focused on studying properties of parameterised curves of equilibria. The end result of performing Lyapunov-Schmidt reduction is a low-dimensional parameterised algebraic system of equations whose zeros are in one-to-one correspondence with equilibria of the original dynamical system near its bifurcation. This allows for local bifurcation analysis without the need to derive explicit approximations of a nonlinear manifold, for which tools from singularity theory can be utilised [Golubitsky1985].
A notable limitation of Lyapunov-Schmidt reduction is the local nature of its results, which hold on arbitrarily small neighbourhoods of a system’s bifurcation point. This limitation can be traced to a key step in the reduction procedure being an application of the Implicit Function Theorem (ImFT), whose conclusions are local. Locality is often not an issue when one’s goal is to simply classify the nature of a bifurcation. However, for many engineering applications it is also important to establish explicit bounds of validity for the analysis, for example in order to quantify robustness of the established features to parameter uncertainty and state perturbation. To the authors’ best knowledge, such bounds have not been formally established in existing literature. In this paper we address this gap and derive estimates for the bounds of validity of Lyapunov-Schmidt reduction, drawing on recent advances in estimating and applying bounds for the domain of validity for the ImFT [jindal2023imft, jindal2023fdbklin].
The primary contributions of the presented work are 1) a derivation of bounds of validity for the Lyapunov-Schmidt reduction procedure for finite-dimensional systems of nonlinear algebraic equations, and 2) detailed computation of these bounds for a low-dimensional illustrative example.
The paper is organised the following way. In section II we introduce notation and background information on the Implicit Function Theorem and the Lyapunov-Schmidt reduction procedure. Then in section LABEL:sec:bounds we establish our main result and derive the bounds. We apply these bounds to a low-dimensional illustrative example in section LABEL:sec:example and conclude in section LABEL:sec:conclusion.
II Preliminaries
II-A Notation and definitions
Let $\mathbb{R}$ denote the set of real numbers and $\mathbb{N}$ denote the set of positive integers. Throughout the text we let $U\subseteq\mathbb{R}^{n}$ and $V\subseteq\mathbb{R}^{m}$ denote open subsets of Euclidean spaces with appropriate dimensions $m,n$. We will say a map $\mathbf{f}:U\to V$ belongs to class $C^{\nu}$ for $\nu\in\mathbb{N}$ if it is $\nu$-times continuously differentiable. An open ball of radius $r>0$ centred at $\mathbf{x}_{0}\in\mathbb{R}^{n}$ is denoted by $\mathcal{B}(\mathbf{x}_{0},r)=\{\mathbf{x}\in\mathbb{R}^{n}:\norm{\mathbf{x}-\mathbf{x}_{0}}<r\}$ where $\norm{\cdot}$ is a vector norm on $\mathbb{R}^{n}$. For any matrix $A\in\mathbb{R}^{s\times t}$ we use $\norm{A}$ to denote a matrix norm induced by the same vector norm, i.e., $\norm{A}:=\sup\{\norm{A\mathbf{x}}:\norm{\mathbf{x}}=1\}$. For a $f\in\mathcal{C}^{\nu}$, $Df$ denotes the Jacobian matrix and $\mathrm{D}_{\mathbf{y}}f$ denotes a Jacobian matrix of partial derivatives with respect to the variables $\mathbf{y}$. $\mathcal{O}(\mathbf{x})$ denotes a neighbourhood of $\mathbf{x}$.
II-B Implicit Function Theorem
Consider a $\mathcal{C}^{\nu}$ map $f:U\times V\to\mathbb{R}^{m}$ with $\nu\geq 1$. The implicit function theorem is a fundamental result in mathematical analysis that establishes conditions under which the variables $\mathbf{x}\in U$ implicitly define the variables $\mathbf{y}\in V$ through the relation $f(\mathbf{x},\mathbf{y})=w_{0}$ for $w_{0}\in\mathbb{R}$ [krantz2002implicit]. We first state a version of this standard result from [zeidler2013vol1, Theorem 4.B].
Theorem II.1 (Implicit Function Theorem)
Let $U\subset\mathbb{R}^{n}$ and $V\subset\mathbb{R}^{m}$ be be open subsets. Consider $\mathcal{C}^{\nu}$ smooth map
$U\times V\ni(\mathbf{x},\mathbf{y})\to f(\mathbf{x},\mathbf{y})\in\mathbb{R}^{m}$ (2.1)
where $\nu\geq 1$. Consider a point $(\mathbf{x}_{0},\mathbf{y}_{0})\in U\times V$ for which $\mathrm{D}_{\mathbf{y}}f(\mathbf{x}_{0},\mathbf{y}_{0}):\mathbb{R}^{m}\to\mathbb{R}^{m}$ is an isomorphism. Then for any neighbourhood $\mathcal{O}(\mathbf{y}_{0})$ of $\mathbf{y}_{0}$ there exists a neighbourhood $\mathcal{O}(\mathbf{x}_{0})$ of $\mathbf{x}_{0}$ and a $\mathcal{C}^{\nu}$ map $g:\mathcal{O}(\mathbf{x}_{0})\to\mathcal{O}(\mathbf{y}_{0})$ satisfying
$f(\mathbf{x},g(\mathbf{x}))=f(\mathbf{x_{0}},\mathbf{y}_{0})\qquad\forall\,\mathbf{x}\in\mathcal{O}(\mathbf{x}_{0})$ (2.2)
In [jindal2023imft, Corollary 3.8], explicit bounds on the regions in which the results of the implicit function hold were presented. We restate this result below.